OpenStudy (anonymous):

For a continous function f(x), f(1) = 11 and integral f ' (x)dx = 9 from 1 to 5. Determine f(5)

OpenStudy (anonymous):

@Nurali

ganeshie8 (ganeshie8):

By FTC : \[f(x) = \int f'(x) dx\]

ganeshie8 (ganeshie8):

\[ \int \limits_1^5 f'(x) dx = 9 \] \[\implies f(5) - f(1) = 9 \] \[\implies f(5) - 11 = 9 \implies f(5) = 20 \]